OpenStudy (anonymous):

Consider a triangle ABC with interior angles A, B and C, and the corresponding opposite sides a,b and c. Show that the area of ABC is given by : [a^(2)sinBsinC]/(2sinA) [URGENT]

OpenStudy (e.mccormick):

OK. So trig ratios.

OpenStudy (e.mccormick):

Now, we know it is a triangle, but not if it is right or not. So we have to use something other than the typical 1/2 bh.

OpenStudy (anonymous):

yeah

OpenStudy (e.mccormick):

We have to arrive at:\[\frac{a^{2}\sin B\sin C}{2\sin A)\]

OpenStudy (anonymous):

yup...

OpenStudy (e.mccormick):

Hmm... the math formatting did not work right for me there. Odd. AH! I see the mistake. \[\frac{a^{2}\sin B\sin C}{2\sin A}\] OK. There we are. If you think about it, there are right triangles hidden in there that tell you a bit about the relationships between the different angles.

OpenStudy (anonymous):

what do you mean?

OpenStudy (e.mccormick):